OpenStudy (anonymous):

what is the factorization of 49x^2+70x+25

OpenStudy (anonymous):

does it help to know that it is a perfect square ?

OpenStudy (anonymous):

\(49x^2=(7x)^2, 25=5^2\) and \(2\times 7\times 5=70\)

OpenStudy (anonymous):

would it (2x+7)(x+5)

OpenStudy (anonymous):

no it is a perfect square

OpenStudy (anonymous):

\((7x+5)^2\) in other words